# Thinking beyond the anthropomorphic paradigm benefits LLM research

### v1

@@ Abstract
Anthropomorphism, or the attribution of human traits to technology, is an automatic and unconscious response that occurs even in those with advanced technical expertise. In this position paper, we analyze hundreds of thousands of computer science research articles from the past decade and present empirical evidence of the prevalence and growth of anthropomorphic terminology in research on large language models (LLMs). This terminology reflects deeper anthropomorphic conceptualizations which shape how we think about and conduct LLM research. We argue these conceptualizations may be limiting, and that challenging them opens up new pathways for understanding and improving LLMs beyond human analogies. To illustrate this, we identify and analyze five core anthropomorphic assumptions shaping prominent methodologies across the LLM development lifecycle, from the assumption that models must use natural language for reasoning tasks to the assumption that model capabilities should be evaluated through human-centric benchmarks. For each assumption, we demonstrate how non-anthropomorphic alternatives can open new directions for research and development.
Machine Learning, ICML
@@ 1 Introduction
When a large language model (LLM) outputs a factually incorrect answer with seeming confidence, we call it a “hallucination.” When it generates inconsistent answers, we describe it as “confused.” These descriptions reflect the deeply-ingrained human tendency to anthropomorphize, or attribute human characteristics to, non-human entities (Epley, 2018). Since anthropomorphism tends to be an automatic, unconscious response (Dacey, 2017), many are not aware of how prevalent it is – or how limiting it can be. Dijkstra (1985) famously wrote that anthropomorphic concepts are “so pervasive that many of my colleagues don’t realize how pernicious it is.” Forty years later, as artificial intelligence (AI) systems become increasingly advanced and prevalent, carefully examining the role of anthropomorphic thinking in LLM research has become essential.
Anthropomorphic conceptualizations play a complex and often productive role in LLM research. For example, chain-of-thought prompting and attention mechanisms, inspired by human cognitive processes, have led to significant advances in the field. Such thinking is inherently imprecise in its application to artificial systems. However, we do not dismiss it on these grounds – nor do we argue against it in general. Instead, this position paper argues that advancing LLM research requires moving beyond our default reliance on anthropomorphic thinking.
Recent scholarship has critically examined the pervasive use of anthropomorphic terminology LLM research (Cheng et al., 2024; Shanahan, 2024). Our quantitative analysis reveals a notable increase in anthropomorphic language over recent years. As linguistic philosophers argue, terminology fundamentally shapes our conceptualization of abstract phenomena (Lakoff & Johnson, 2008). For instance, labeling an AI error as a “hallucination” subtly influences perceptions of its origin and potential remediation. While anthropomorphic terms are readily observable, they represent only the tip of the iceberg of anthropomorphic conceptualizations. This paper seeks to unpack not just the visible anthropomorphic terminology, but also the implicit assumptions and methodological frameworks that underlie these linguistic choices.
@@
To do so, we present a framework for analyzing how anthropomorphic assumptions can shape and actively limit research directions, and we identify alternatives that challenge these assumptions (Figure 1). We apply our framework to analyze five assumptions across the LLM development and deployment lifecycle. For model pretraining, we identify and challenge the assumption that human-like methods are optimal for enabling models to perform tasks, and instead highlight non-anthropomorphic approaches like tokenizing language with bytes rather than human-understandable words. In alignment, we question the assumption that models must explicitly reason about and implement human values to benefit humanity and instead demonstrate how to leverage model properties that lack human analogs. For measurement and evaluation, we unify critiques of the widespread reliance on human-centric benchmarks to assess model capabilities. For understanding model behavior, we address the assumption that human-like normative judgments or intentions should be assigned to model behaviors. Finally, in end-user interactions, we challenge the notion that human-AI interaction mirrors human-to-human communication. These examples illustrate how pervasive anthropomorphic conceptualizations shape LLM research, and demonstrate the value of thinking beyond them.
@@
Our work contributes the following:
  1. 1.
    Quantitative evidence of anthropomorphic terminology’s growing prevalence in computer science, particularly in LLM research, through analysis of over 250,000 research abstracts.
  2. 2.
    A framework for examining how anthropomorphic assumptions shape LLM research methodology across five key stages of development and deployment.
  3. 3.
    Concrete research directions that complement existing approaches by moving beyond anthropomorphic assumptions.
Figure 1: A framework for analyzing anthropomorphic conceptualizations across the LLM development and deployment lifecycle. Anthropomorphism affects not only the terminology we use, but also the conceptualizations we have (Medin, 1989; Murphy, 2004) and, in turn, the research questions that we pursue and methodologies that we develop.
@@ 2 Background & related work
2.1 Anthropomorphism & human-AI interaction
@@
Research on anthropomorphic perceptions of computer systems spans fields like human-computer interaction, psychology, and cognitive science. Work with ELIZA, an early NLP program from the 1960s, demonstrated how even simple pattern-matching can elicit strong anthropomorphic responses from users, who attributed understanding and empathy to the program (Weizenbaum, 1966). The Computers As Social Actors paradigm later established that humans inherently apply social expectations to computers (Nass et al., 1994). More recently, research has identified various factors that increase anthropomorphic perceptions, from visual features in embodied agents to linguistic cues like expressions of emotions (Glaese et al., 2022; Abercrombie et al., 2023; DeVrio et al., 2025). Studies of commercial AI assistants like Siri and Alexa have further documented how interface elements like personalized names and simulated personality traits shape users’ mental models of what AI can do (Abercrombie et al., 2021). Many user studies demonstrate that anthropomorphic perceptions impact human-AI interaction, shaping trust, reliance, sensitive disclosures, and emotional attachment (Li et al., 2024; Song & Luximon, 2020; Zhou et al., 2024; Khadpe et al., 2020; Bender, 2024; Abercrombie et al., 2023; Mozafari et al., 2020; Gros et al., 2021). In particular, scholars have associated anthropomorphic perceptions with facilitating inappropriate levels of trust in AI and inflating expectations of what AI can do (Winkle et al., 2021; Inie et al., 2024). Many of these effects have been documented in both technical novices and experts (Nass et al., 1999).
2.2 History of anthropomorphism in AI research
@@
In computer science, there is a long history of borrowing concepts from cybernetics and cognitive science to characterize novel phenomena (Floridi & Nobre, 2024). This is especially prominent in the subfields of artificial intelligence (AI) and machine learning, which from their nascence, aimed to reproduce aspects of human intelligence (Brynjolfsson, 2023). Key events in the nascence of AI are emblematic of how anthropomorphic conceptualizations have shaped this field from its conception. First, Turing (1950)’s article titled “Computing Machinery and Intelligence” introduces an “imitation game”, or what has become better known as the Turing test, to determine if a machine is capable of “intelligence” and “thinking.” Later work has problematized the anthropomorphic attributions of this test, pointing out that it actually tests whether a machine’s output can fool a human into believing the illusion that it can think, rather than measuring real cognitive processes (Proudfoot, 2011). We further unpack this conflation between demonstrated human-like ability and underlying processes in Section 4.3.
The coining of the term “artificial intelligence” is often attributed to the 1956 Dartmouth College workshop organized by John McCarthy and others to discuss research oriented towards solving “problems now reserved for humans” (McCarthy et al., 1956, 2006). Across these events, research ideas are consistently framed using anthropomorphic terminology, inextricable from the anthropomorphic concepts on which the field continues to rely. While these questions about replicating human-like abilities remain powerful and exciting decades later, we argue that they can limit language model development, and that non-anthropomorphic thinking can unlock new avenues of progress.
2.3 Critiques of anthropomorphism
@@
Our work builds on existing critiques of anthropomorphism in (computer) science. Shanahan (2024) caution against anthropomorphic language when describing language models, arguing for more technical precision and new metaphors. Dai (2024) argue that a “mechanistic view” of AI, which implicitly treats AI as a human-like agent capable of moral decision-making, is a flawed approach that ultimately hinders establishing accountability for AI harms. These more recent works build on decades of critique, tracing back to as early as Dijkstra (1985), who disparages the prevalent use of anthropomorphic terminology in science more broadly, arguing that it is more misleading than helpful because we lose control over the human-like connotations associated with this terminology. We expand on this point and concretize it by not only examining terminology but also surfacing the assumptions and methodologies that come with them.
@@ 3 Prevalence of anthropomorphism in recent research on LLMs
Previous work has shown that anthropomorphism is prominent and rapidly increasing in computer science research over the past decades, with papers on natural language processing (NLP) and LLMs exhibit the highest levels of anthropomorphic framing (Cheng et al., 2024). We quantitatively demonstrate that this prevalence has only increased in recent years. Cheng et al. (2024) quantify this using AnthroScore, a measure of implicit anthropomorphic framing in language used to describe technologies. AnthroScore uses the masked language model RoBERTa to calculate the relative probability that a given entity $x$ (e.g., “language model”) in a sentence $s$ would be appropriately replaced by human pronouns (“he”, “she”) versus non-human pronouns (“it”). Specifically, the degree of anthropomorphism for entity $x$ in sentence $s$ is measured as
$A(s_{x})=\log\frac{P_{\textsc{human}}(s_{x})}{P_{\textsc{non-human}}(s_{x})}$ (1)
where
$P_{\textsc{human}}(s_{x})=\sum_{w\in\text{human pronouns}}P(w),$
$P_{\textsc{non-human}}(s_{x})=\sum_{w\in\text{non-human pronouns}}P(w),$
@@
We compute $A^{\text{bin}}(S)$ on a dataset of over 200,000 computer science papers posted on arXiv from January 2023 – December 2024 (the most recent data available from arXiv.org submitters (2024)) that mention a “system”, “network”, or “model” (following the approach of Cheng et al. (2024)). The longitudinal trend is presented in Figure 2. We find that anthropomorphism is generally prevalent, with 34% of abstracts having anthropomorphism in January 2023, and this number steadily increasing to 40% by December 2024. (For each abstract, we define having anthropomorphism as $A^{\text{bin}}(S)=1$.) More strikingly, for papers mentioning LLMs [^1] We define this following the method of Movva et al. (2024) as papers mentioning terms such as “large language model”, “foundation model”, “llama”, “gpt”, etc., over 40% of abstracts have anthropomorphism in January 2023, and this number also rises to 48% by December 2024. This reveals both the prevalence and growing use of anthropomorphic framing in computer science research and particularly in LLM research.
@@
We also compute this new metric $A^{\text{bin}}(S)$ on the $>50,000$ abstracts in the ACL Anthology dataset from 2007 - 2022 to reproduce the findings from Cheng et al. (2024), but aggregating over the abstracts using $A^{\text{bin}}(S)$ rather than on the sentence level (Figure 3). Corroborating their finding of a steady increase, we find that the percentage of anthropomorphic abstracts has more than doubled, increasing from 5% to 11%.
@@
In the ACL anthology, we find significant differences in anthropomorphism across NLP subfields. Using the model-predicted topic labels provided by the ACL anthology, we compare $A^{\text{bin}}(S)$ across different topics (Figure 4). We find that the categories of “Interpretability and Analysis of Models for NLP”, “Ethics and NLP”, and “Dialogue and Interactive Systems” have the highest percentages of anthropomorphic abstracts. This trend aligns with these fields’ recent surge in popularity and their increasing focus on LLMs. Anthropomorphic assumptions are particularly embedded in model analysis, ethical questions, and user-facing interactive systems. We unpack the impact and limitations of these assumptions in Section 4, which especially motivates our discussions in Sections 4.2, 4.4, and 4.5. Our finding of ethics having high rates of anthropomorphism also builds on previous work problematizing assumptions of agenthood in ethics analyses (Dai, 2024). In contrast, more classical subfields of NLP that do not involve LLMs, such as discourse and pragmatics, syntax, and semantics have the lowest rates of anthropomorphism.
Figure 2: Temporal increase in anthropomorphic abstracts in computer science arXiv papers mentioning “models” or “systems” from January 2023 – October 2024. Shading reflects 95% CI. Based on the percent of abstracts with at least one anthropomorphic sentence, we find that anthropomorphism is prevalent and is steadily increasing, especially in LLM papers.
Figure 3: Temporal increase in anthropomorphic abstracts in ACL Anthology papers from 2007 – 2022. Shading reflect 95% CI. Based on the percent of abstracts with at least one anthropomorphic sentence, we find that anthropomorphism has doubled in the past decade.
Figure 4: Rates of anthropomorphic abstracts by ACL anthology topics. Error bars reflect 95% CI. “Interpretability”, “ethics”, and “dialogue” have the highest rates of anthropomorphism, reflecting the prevalence of anthropomorphic assumptions in these areas, which we explore in Section 4.

### v2

@@ Abstract
Anthropomorphism, or the attribution of human traits to technology, is an automatic and unconscious response that occurs even in those with advanced technical expertise. In this position paper, we analyze hundreds of thousands of research articles to present empirical evidence of the prevalence and growth of anthropomorphic terminology in research on large language models (LLMs). We argue for challenging the deeper assumptions reflected in this terminology — which, though often useful, may inadvertently constrain LLM development — and broadening beyond them to open new pathways for understanding and improving LLMs. Specifically, we identify and examine five anthropomorphic assumptions that shape research across the LLM development lifecycle. For each assumption (e.g., that LLMs must use natural language for reasoning, or that they should be evaluated on benchmarks originally meant for humans), we demonstrate empirical, non-anthropomorphic alternatives that remain under-explored yet offer promising directions for LLM research and development.
@@ 1 Introduction
When a large language model (LLM) outputs a factually incorrect answer with seeming confidence, we call it a “hallucination.” When it generates inconsistent answers, we describe it as “confused.” These descriptions reflect the deeply-ingrained tendency to anthropomorphize, or attribute human characteristics to, non-human entities [29]. As anthropomorphism tends to be an automatic, unconscious response [23], many are not aware of its prevalence or influence. This has contributed to anthropomorphism in science being a recurring and long-standing issue; in 1985, Dijkstra [27] famously wrote that it is “so pervasive that many of my colleagues don’t realize how pernicious it is.”
Forty years later, anthropomorphism plays a complex role in LLM research. It offers intuitive scaffolding for complex concepts and allows researchers to draw on insights from cognitive science and psychology, which has enabled significant advances like instruction-tuning and chain-of-thought prompting. In this position paper, we argue that while anthropomorphism can be productive, the field has become overly reliant on it; advancing LLM research requires moving beyond our default dependence on anthropomorphic thinking. We advocate for a both-and, rather than an either-or approach, calling for recognizing not only when anthropomorphic thinking is useful, but also when it constrains research and limits the discovery of novel capabilities.
To do so, we first demonstrate the prevalence of anthropomorphism, beginning with a quantitative analysis of recent LLM research articles. This reveals a notable increase in anthropomorphic terminology over the years ($\sim 150\%$ increase since 2007). Moreover, while anthropomorphic terms are easiest to measure, they represent only the tip of the iceberg of anthropomorphic thinking. We unpack not just the visible anthropomorphic terminology, but also the implicit assumptions that underlie these linguistic choices and shape the resulting research. We present a framework for analyzing how anthropomorphic assumptions have shaped, but may also limit research directions (Table 1).
@@
We apply our framework to analyze five assumptions across the LLM development and deployment lifecycle, unifying recent empirical results and discourse to highlight the value of under-explored, non-anthropomorphic approaches. For model pretraining, we identify and challenge the assumption that human-like methods are optimal for enabling models to perform tasks, and instead highlight non-anthropomorphic approaches like tokenizing language with bytes rather than human-understandable words. In alignment, we question the assumption that models must explicitly reason about and implement human values to benefit humanity and instead demonstrate how to leverage model properties that lack human analogs. For measurement and evaluation, we unify critiques of the widespread reliance on human-centric benchmarks to assess model capabilities. For understanding model behavior, we address the assumption that human-like normative judgments or intentions should be assigned to model behaviors. Finally, in end-user interactions, we challenge the notion that human-AI interaction mirrors human-to-human communication. These examples highlight the pervasiveness of anthropomorphism and emphasize the potential of non-anthropomorphic solutions.
@@
Our work contributes the following: (1) a large-scale quantitative analysis of over 250,000 research abstracts, revealing the increasing prevalence of anthropomorphic terminology in computer science, especially in LLM research; (2) a framework for analyzing how anthropomorphic assumptions influence LLM research across five key stages of development and deployment; and (3) a set of complementary research directions that challenge each anthropomorphic assumption, unified to highlight the value of non-anthropomorphic approaches.
Table 1: Summary of our framework to analyze the tradeoffs of anthropomorphism, with examples across the LLM lifecycle. Anthropomorphism affects not only our terminology, but also our assumptions [61, 68] and, in turn, our research questions and methodologies.
Stage Anthropomorphic assumption Examples of how anthropomorphism has been useful Examples of non-anthropomorphic paths forward
Training Human-like approaches are optimal for models. Subword tokenization, chain-of-thought reasoning Byte-level tokenization, solving reasoning tasks in models’ latent space
Alignment Models should explicitly reason about and implement human values to be safe & helpful. Reinforcement learning from human feedback, constitutional AI, instruction-tuning Developing normative specifications without morals, drawing from control systems theory, steering models using mechanistic interpretability
Evaluation Model capabilities should be measured in human-like ways. Using existing standardized/multiple choice tests for evaluation, static behavioral benchmarks Dynamic evaluations that reflect human-LLM interaction, designing tests which take into account model-specific challenges and phenomena
Understanding model behavior Human-like normative judgments or intentions should be assigned to human-like model behaviors. Characterizing phenomena like hallucinations, sycophancy, and deception Refraining from assigning normative value or intention to LLM outputs; understanding LLM behaviors as simulations rather than reflecting internal states
User interaction Human-LLM interactions mirror human-human interactions. Prompting and user-friendly conversational interfaces Structured input formats to improve LLM performance, interfaces that more accurately reflect system capabilities
@@ 2 Background & related work
Anthropomorphism & human-AI interaction
@@
Research on anthropomorphic perceptions of computer systems spans fields like human-computer interaction, psychology, and cognitive science. Work with ELIZA, an early NLP program from the 1960s, demonstrated how even simple pattern-matching can elicit strong anthropomorphic responses from users, who attributed understanding and empathy to the program [87]. The Computers As Social Actors paradigm later established that humans inherently apply social expectations to computers [69]. More recently, research has identified various factors that increase anthropomorphic perceptions, from visual features in embodied agents to linguistic cues like expressions of emotions [34, 1, 26, 41]. Many user studies demonstrate that anthropomorphic perceptions impact human-AI interaction, shaping trust, reliance, sensitive disclosures, and emotional attachment [55, 81, 93, 48, 12, 1, 67, 36]. In particular, scholars have associated anthropomorphic perceptions with facilitating inappropriate levels of trust in AI and inflating expectations of what AI can do [88, 42]. Many of these effects have been documented in both technical novices and experts [70].
History of anthropomorphism in AI research
@@
In computer science, there is a long history of borrowing concepts from cybernetics and cognitive science to characterize novel phenomena [32]. This is especially prominent in the subfields of artificial intelligence (AI) and machine learning, which from their nascence, aimed to reproduce aspects of human intelligence [16]. Key events in the nascence of AI are emblematic of how anthropomorphic thinking has shaped this field from its conception. First, Turing [84] introduces an “imitation game”, or what has become better known as the Turing test, to determine if a machine is capable of “intelligence” and “thinking.” Later work problematized the anthropomorphic attributions of this test, pointing out that it actually tests whether a machine’s output can fool a human into believing the illusion that it can think, rather than measuring real cognitive processes [75]. The coining of the term “artificial intelligence” is often attributed to a 1956 workshop organized to discuss research oriented towards solving “problems now reserved for humans” [57, 58]. Across these events, research ideas are consistently framed using anthropomorphic terminology, inextricable from the anthropomorphic concepts on which the field continues to rely. While these questions about replicating human-like abilities remain powerful and exciting decades later, we argue that thinking beyond them in non-anthropomorphic ways can further unlock new avenues of progress.
Critiques of anthropomorphism
@@
Our work builds on existing critiques of anthropomorphism in (computer) science. Shanahan [78] caution against anthropomorphic language when describing language models, arguing for more technical precision and new metaphors. Dai [24] argue that a “mechanistic view” of AI, which implicitly treats AI as a human-like agent capable of moral decision-making, is a flawed approach that ultimately hinders establishing accountability for AI harms. These more recent works build on decades of critique, tracing back to as early as Dijkstra [27], who disparages the prevalent use of anthropomorphic terminology in science more broadly, arguing that it is more misleading than helpful because we lose control over the human-like connotations associated with this terminology. We expand on this point and concretize it by not only examining terminology but also surfacing the impacts of the assumptions that come with them.
@@ 3 Prevalence of anthropomorphism in recent research on LLMs
Previous work has shown that anthropomorphism is prominent and rapidly increasing in computer science research over the past decades, with papers on natural language processing (NLP) and LLMs exhibit the highest levels of anthropomorphic framing [20]. We quantitatively demonstrate that this prevalence has only increased in recent years. Cheng et al. [20] quantify this using AnthroScore, a measure of implicit anthropomorphic framing in language used to describe technologies. AnthroScore uses the masked language model RoBERTa to calculate the relative probability that a given entity $x$ (e.g., “language model”) in a sentence $s$ would be appropriately replaced by human pronouns (“he”, “she”) versus non-human pronouns (“it”). Specifically, the degree of anthropomorphism for entity $x$ in sentence $s$ is measured as
$A(s_{x})=\log\frac{P_{\textsc{human}}(s_{x})}{P_{\textsc{non-human}}(s_{x})},P_{\textsc{human}}(s_{x})=\sum_{w\in\text{\{he, she\}}}P(w),P_{\textsc{non-human}}(s_{x})=\sum_{w\in\text{\{it\}}}P(w),$
@@
We compute $A^{\text{bin}}(S)$ on a dataset of over 200,000 computer science papers posted on arXiv from January 2023 – December 2024 (the most recent data available from arXiv.org submitters [7]) that mention a “system”, “network”, or “model” (following the approach of Cheng et al. [20]). The longitudinal trend is presented in Figure 1 (left). Anthropomorphism is generally prevalent, with 34% of abstracts having anthropomorphism in January 2023, and this number steadily increasing to 40% by December 2024. (For each abstract, we define having anthropomorphism as $A^{\text{bin}}(S)=1$.) More strikingly, for papers mentioning LLMs [^1] We define this following the method of Movva et al. [66] as papers mentioning terms such as “large language model”, “foundation model”, “llama”, “gpt”, etc., over 40% of abstracts have anthropomorphism in January 2023, and this number also rises to 48% by December 2024. This reveals both the prevalence and growing use of anthropomorphic framing in computer science (and especially LLM) research.
@@
We also compute this new metric $A^{\text{bin}}(S)$ on the $>50,000$ abstracts in the ACL Anthology dataset from 2007 - 2022 to reproduce the findings from Cheng et al. [20], but aggregating over the abstracts using $A^{\text{bin}}(S)$ rather than on the sentence level (Figure 1, middle). Corroborating their finding of a steady increase, we find that the percentage of anthropomorphic abstracts has more than doubled, increasing from 5% to 11%.
@@
In the ACL anthology, we find significant differences in anthropomorphism across NLP subfields. Using the model-predicted topic labels provided by the ACL anthology, we compare $A^{\text{bin}}(S)$ across different topics (Figure 1). We find that the categories of “Interpretability and Analysis of Models for NLP”, “Ethics and NLP”, and “Dialogue and Interactive Systems” have the highest percentages of anthropomorphic abstracts. This trend aligns with these fields’ recent surge in popularity and their increasing focus on LLMs. Anthropomorphic assumptions are particularly embedded in model analysis, ethical questions, and user-facing interactive systems. We unpack the impact and limitations of these assumptions in Section 4, which especially motivates our discussions in Sections 4.2, 4.4, and 4.5. Our finding of ethics having high rates of anthropomorphism also builds on previous work problematizing assumptions of agenthood in ethics analyses [24]. In contrast, more classical subfields of NLP that do not involve LLMs, such as discourse and pragmatics, syntax, and semantics have the lowest rates of anthropomorphism.
Figure 1: Temporal increase in % of abstracts with $>1$ anthropomorphic sentence in CS arXiv papers from Jan 2023 – Oct 2024 (left) and ACL anthology papers from 2007 – 2022 (middle). Anthropomorphism is prevalent and is steadily increasing, especially in LLM and NLP papers. Rates of anthropomorphic abstracts by ACL anthology topics (right). “Interpretability", “ethics", and “dialogue" have the highest rates of anthropomorphism, reflecting the prevalence of anthropomorphic assumptions in these areas, which we explore in Section 4. Shading and error bars reflect 95% CI.
Score sensitivity
To understand the cause of the prevalence of anthropomorphism, we perform a Fightin’ Words analysis [64] on our dataset (following the approach of Cheng et al. [20]), identifying verbs with highest z-scores for high ($>1$) and low ($<-1$) AnthroScore sentences, i.e., the words that differ most significantly in distribution between these two sets using weighted log odds ratios. This also enables us to understand the sensitivities and biases of the AnthroScore metric. The results are as follows:
High AnthroScore: achieve, guide, demonstrate, teach, ask, train, prompt, follow, make, target, become, learn, understand, excel, require, mislead, answer, hallucinate, memorize, draw;
Low AnthroScore: propose, outperform, use, develop, present, evaluate, enhance, improve, introduce, implement, validate, apply, reduce, adapt, employ, extend, allow, leverage, utilize, design.
Top verbs for high AnthroScore reflect both human-like actions and also verbs that are commonly used in discussing LLMs, underscoring the prevalence of anthropomorphic terminology in papers about LLMs.
Significance
Terminology is an imperfect proxy for anthropomorphic thinking. However, building on prior work demonstrating that language fundamentally structures our thinking [33, 52, 15, 43], it nonetheless is a tractable measurement approach that provides valuable signal about trends in underlying assumptions. In the next section, we turn to these assumptions in more detail.